If n=10, ¯ x x ¯ (x-bar)=49, and s=6, find the margin of error at a 98% confidence level Give your answer to two decimal places.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 49

sample standard deviation = s = 6

sample size = n = 10

Degrees of freedom = df = n - 1 = 10 - 1 = 9

At 98% confidence level

\alpha = 1 - 98%

\alpha =1 - 0.98 =0.02

\alpha/2 = 0.01

t\alpha/2,df = t0.01,9 = 2.821

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.821 * (6 / \sqrt 10)

Margin of error = E = 5.35